OpenStudy (anonymous):

How do I find the equation of this graph? I'm guessing that this formula: y=a√(b(x-h))+k has something to do with it. Thank you :)

OpenStudy (anonymous):

Here's the graph :P

OpenStudy (jdoe0001):

http://hotmath.com/hotmath_help/topics/graphing-square-root-functions/graph-1.gif yours is shifted 3 units to the left A( Bx + C ) + D C = horizontal shift, C < 0, to the right, C>0, to the left

OpenStudy (anonymous):

@jdoe0001 Okay, but how do you find all the numbers to get the answer? Do you just have to compare to y = √x or something?

OpenStudy (jdoe0001):

yep, you just compare the derived function from the parent function so yours is shifted to the LEFT by 3 units so it will be \(y=\sqrt{x+3}\)

OpenStudy (jdoe0001):

hmm, yours seem to be running a little higher on the "y" axis, so it may have a multiplier too

OpenStudy (jdoe0001):

A( Bx + C ) + D A = shrinks or expands the graph horizontally

OpenStudy (anonymous):

Oh okay :) and yeah i agree with what you said for A

OpenStudy (anonymous):

Whoops, sorry I just re-looked at my information and for A, it's a vertical stretch and a < 0 means is reflected across the x-axis

OpenStudy (jdoe0001):

that's correct, if A is negative, the graph goes upside down

OpenStudy (jdoe0001):

so-called, reflected over x-axis